We are studying the lesson, Adding Real Numbers this week in class.
Can you help me with the problem variable x divided by 4 plus negative variable x divided by three? in numeral form x/4+-x/3=???
Thank you!

Answer by Alan3354(30993)

x/4+-x/3=???
= x/4 - x/3
= 3x/12 - 4x/12
= -x/12